Given the points (-2,2) and (1,3) on a line, calculate the slope of that line
Please help!!!!!!!!!!
Lackinmath

Answer by uma(370)

Slope = difference in y co ordinates/ difference in x co ordinates
==> slope = (3 - 2)/(1 - (-2))
= 1/1+2
==> slope = 1/3


Good Luck!!!